OpenStudy (czesc):

INSTANT FAN MEDAL OR SNAPCHAT ;) Find the corresponding rectangular equation represented by the parametric equations x = 1 + sec θ and y = 1 + tan θ by eliminating the parameter.

OpenStudy (czesc):

I solved for theta obviously, but I'm not sure how to simplify that further. Is there an identity I should be using?

OpenStudy (czesc):

these are the options http://puu.sh/oGsbP/b09533b4d2.png

OpenStudy (nfcfox):

\[x=1+\sec \theta--->\sec \theta=x-1--->\frac{ 1 }{ \cos \theta}=x-1--->\cos \theta=\frac{ 1 }{ x-1}\]

OpenStudy (nfcfox):

\[y=1+\tan \theta--->y=1+\frac{ \sin \theta }{ \cos \theta}--->(y-1)\cos \theta=\sin\]

OpenStudy (nfcfox):

Plug costheta into that equation

OpenStudy (freckles):

I would square one of the equations given (you use Pythagorean identity ) it should be easy to write in terms of x and y then

OpenStudy (czesc):

@freckles @nfcfox that makes it a lot simpler, thank you guys so much, i will update further with solution

OpenStudy (kawii2004):

are u k12

OpenStudy (nfcfox):

\[\sin^2\theta+\cos^2\theta=1---> ((y-1)(\frac{ 1 }{ x-1}))^2+(1/(x-1))^2=1\]

OpenStudy (nfcfox):

@czesc what's your sc

OpenStudy (czesc):

@nfcfox once i finish this assignment I'll send it to you privately so i dont get a billion adds

OpenStudy (nfcfox):

@czesc ok

OpenStudy (czesc):

btw what'd you get for the final rectangular equation @nfcfox

OpenStudy (czesc):

in terms of these options: http://puu.sh/oGsbP/b09533b4d2.png @nfcfox because I can't get anything like that lol

OpenStudy (freckles):

hey @czesc this is what I did \[x=1+\sec(\theta) \\ \text{ square both sides } \\ x^2=(1+\sec(\theta))^2 \\ x^2=1+2 \sec(\theta)+\sec^2(\theta) \\ x^2=1+2 \sec(\theta)+1+\tan^2(\theta) \text{ by pythagorean identity } \\ x^2=2 +2 \sec(\theta)+\tan^2(\theta) \text{ adding like terms } \\ \\ \text{ now we can use our given equations to write } \\ \text{ the right hand side in terms of } x \text{ and } y\]

OpenStudy (freckles):

\[x=1+ \sec(\theta) \implies x-1=\sec(\theta) \\ y=1+\tan(\theta) \implies y-1=\tan(\theta)\]

OpenStudy (freckles):

plug in

OpenStudy (czesc):

@freckles that makes a lot of sense now, thank you so much

OpenStudy (freckles):

you will have to expand a little and get an equation equal to 0 to compare to your choices but you shouldn't have too much more work to do from that above :)

OpenStudy (czesc):

@freckles yep you pretty much solved it lol

OpenStudy (czesc):

@freckles i wish i could give more than one medal :)

OpenStudy (freckles):

you can give the medal to the other guy i have cookies at home that I can reward myself with

OpenStudy (czesc):

haha see you around thanks again!

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

if it was me, i would cheat and get it quickly maybe

OpenStudy (misty1212):

if you put \(\theta)=0\) you have the point \((2,1)\) on the graph since you have choices, you can check